OpenStudy (anonymous):

what is the full expansion of (3x+2)^4

OpenStudy (bahrom7893):

rewrite as: [3x+2]^2 * [3x+2]^2

OpenStudy (bahrom7893):

(9x^2 + 2* 6x + 4) * (9x^2 + 2* 6x + 4)

OpenStudy (bahrom7893):

(9x^2 + 12x + 4) * (9x^2 +12x +4)

OpenStudy (bahrom7893):

now multiply these out

OpenStudy (bahrom7893):

still workin on it.. have to write it out

OpenStudy (anonymous):

is it 192x?

OpenStudy (bahrom7893):

cant be

OpenStudy (anonymous):

isnt there a chart ypu can use?

OpenStudy (bahrom7893):

no im doin this by hand and calculator

OpenStudy (anonymous):

ithink there is a chart first column is combinations second is the first term third column is the 2nd term...sound familiar?

OpenStudy (bahrom7893):

You mean Pascal's triangle? Sorry I can figure out how to use it, but it will take a while, havent used it in a long time

OpenStudy (bahrom7893):

im just gonna finish off

OpenStudy (anonymous):

yeah using that

OpenStudy (bahrom7893):

Anyway here's the solution: \[81x^4 + 108x^3 + 36x^2 + 108x^3 +144x^2 +48x +36x^2+48x+16 = 81x^4 +216x^3 +216x^2 +96x + 16\]

OpenStudy (bahrom7893):

\[81x^4 + 108x^3 + 36x^2 + 108x^3 +144x^2 +48x +36x^2+48x+16 = 81x^4 +216x^3 +216x^2 +96x + 16\]

OpenStudy (bahrom7893):

dang it the whole thing wont show up brb

OpenStudy (bahrom7893):

81x^4 + 108x^3 + 36x^2 + 108x^3 +144x^2 +48x +36x^2+48x+16 = =81x^4 +216x^3 +216x^2 +96x + 16

OpenStudy (bahrom7893):

that's the final answer, i triple checked my work

OpenStudy (anonymous):

ok thanks